OpenStudy (anonymous):

How can you use the vertex and the sign of "a" to determine the range of the absolute value function? abs() = absolute value sign y = -2abs(x-1)+4

OpenStudy (anonymous):

as i wrote in the last post, \[y=-2|x-1|+4\leq 4\] that is a synonym for saying that the range is \[y\leq 4\]

OpenStudy (anonymous):

you know it is less than or equal to (not greater than) also for the reasons stated in the last post. you know \[|x|\geq 0\] for all x and therefor, because -2 is negative, you know \[-2|x-1|\leq 0\] for all x